X3+x2y-xy2+y3/x2y+xy2

whetther homogeneous or not by dx/dy



1
Expert's answer
2021-05-20T13:57:07-0400

Ans:-

"F(x,y)=\dfrac{x^3+x^2y-xy^2+{y^3}}{{x^2y}+xy^2}\\"


For finding equation homogeneous or not

Put "x=\lambda x" and "y=\lambda y"


"F(\lambda x,\lambda y)=\dfrac{\lambda^3 x^3+\lambda^3 x^2 y-\lambda^3 xy^2+\lambda^3 y^3}{\lambda^3 x^2y+\lambda^3 xy^2}"


"\Rightarrow F(\lambda x,\lambda y)=\dfrac{x^3+x^2y-xy^2+{y^3}}{{x^2y}+xy^2}\\"


"\Rightarrow F(\lambda x,\lambda y)=F(x,y)"


Thus "F(x,y)" is Homogeneous function of degree zero. Therefore, the given differential equation is homogeneous differential equation.